Exhibit 15-6
Below you are given a partial Excel output based on a sample of 16 observations.

ANOVA
df SS MS F
Regression 4,853 2,426.5
Residual 485.3
Coefficients Standard Error
Intercept 12.924 4.425
x1 -3.682 2.630
x2 45.216 12.560


Carry out the test to determine if there is a relationship among the variables at the 5% level. The null hypothesis should ______.

a. be revised
b. not be rejected
c. be rejected
d. None of these answers are correct.
Answer #1

from abvoe test statistic F =2426.5/485.3 =5

for 2,13 degree of freedom, crtiical value at 5 % level =finv(0.05,2,13)=3.806

since test statistic > critical value ; option C is correct

c. be rejected